Warren inveests $4400 in two different bank accounts. This first account pains 12%, the second account paid 11% in interest. At the end of the first year he had earned $513 in intrest. How much was in each account?

Answer by Boreal(15235)

x=first account
4400-x=second account
.12x+.11(4400-x)=513
.12x+484-.11x=513
.01x=29
x=$2900@12%=$348
4400-x=1500@11%=$165
add to $513.